OpenStudy (anonymous):

How do you find the exact value of sin(cot^-1 (5/4))?

OpenStudy (anonymous):

cot^-1(5/4)=x cotx=5/4 cosx/sinx=5/4 cos^2(x)+sin^2(x)=1 sinx=4/sqrt41

OpenStudy (anonymous):

Thank you. Could you explain the part after cosx/sinx=5/4 ?I'm not really following it.

OpenStudy (anonymous):

cotx=cosx/sinx did you know that?

OpenStudy (anonymous):



OpenStudy (anonymous):

I know that cotx=cosx/sinx, I just wasn't really following on what you did after that.

OpenStudy (anonymous):

cos^2 x+sin^2 x=1 cosx=5m and sinx=4m 25m^2+16m^2=1 41m^2=1 m=1/sqrt41 so sinx=4m=4/sqrt41 do you understand now?

OpenStudy (anonymous):

Yes, that makes more sense. Thank you.

OpenStudy (anonymous):

welcome. if i could help, i'm so happy :)